### v1

# Role of Interlayer Shear Phonons on the Lattice Symmetry Switch in a Transition-Metal Dichalcogenide

Abstract
Coherent phonon control using ultrashort pulse trains is the key to realizing structural phase transitions in solids by non-thermal pathways. By combining double-pulse excitation and time-resolved second harmonic generation techniques under high-density electronic excitation in a 2D layered material, WTe2, we demonstrate that the lattice symmetry switching from the Weyl semimetallic Td to the semimetallic 1T′ phases is independent of the amplitude of the coherent interlayer shear phonons. This finding provides new insights into the mechanisms for symmetry switching that electronic excitation-driven shear sliding plays a dominant role.
Photons excite multiple degrees of freedom in matter and induce cooperative phase transitions between electronic and lattice systems via non-thermal pathways. This phenomenon, known as photo-induced phase transition (PIPT) [1], has been a prominent theme in ultrafast spectroscopy applied to topics ranging from photochemistry to solid-state physics for several decades. In classical (incoherent) PIPT, a new phase emerges due to energy relaxation of electronic excited states [2], while in quantum (coherent) PIPT, a transition to a new phase occurs due to coherent control of the electronic or lattice system [3, 4]. Variable time-resolved measurements have made it possible to realize and observe PIPT in liquid crystals [5], phase-change materials [6], and correlated materials [7, 8]. Furthermore, control of ultrafast phase transition based on coherent PIPT is crucial not only for understanding the origin of phase transitions but also for applications such as terahertz frequency-driven optoelectronic memory devices [9]. One promising approach for coherent PIPT is optical pulse-train excitation by tuning the time interval between pump pulses [10, 11, 12]. This excitation scheme allows for the exploration of far non-equilibrium states and coherent control of electronic and structural phases, which is not possible with single-pulse excitation. For example, sub-picosecond electron interference and scattering dynamics [13, 14, 15], photo-induced hidden states [16, 17, 18, 19], and optical control of structural phases [20, 4] have been investigated.
Recently, a feasible system for optical control of coherent PIPT involves ultrafast lattice symmetry switching between the semimetallic 1T′ (centrosymmetric) and Weyl semimetallic Td (non-centrosymmetric) phases of layered transition-metal dichalcogenides (TMDs) such as and [21, 22, 23, 24]. Notably, driving a coherent interlayer shear phonon displacement provides a means to control the lattice symmetry and allow periodic alternation of adjacent structural phases at sub-THz frequencies. To realize coherent PIPT for optical control, it is essential to elucidate the dynamics of incoherent and coherent PIPT in TMDs, focusing on the effect of interlayer shear displacement. Additionally, under high-density electronic excitation conditions exceeding several by a single pulse, the intended phase transition may be suppressed by sample damage or saturable absorption [25, 26, 27]. These issues can arise not only in controlling lattice symmetry but also in realizing optical control of other coherent PIPTs, such as spin states [28], and must be addressed.
In this paper, we address these issues through experimental studies on lattice symmetry change under high-density electronic excitation using time-resolved second harmonic generation (TR-SHG), a sensitive probe for tracking the symmetry change [21, 22, 29, 30, 31]. By combining double-pulse excitation and TR-SHG methods in a two-dimensional layered material system, WTe2, we show that the symmetric switch from the broken-inversion Weyl semimetal Td phase to inversion-symmetric semimetal 1T′ phase is independent of the amplitude of the coherent shear phonons, which has been believed to play a central role.
The sample used in this study was a bulk single crystal of (001) surface Td-WTe2 (from HQ Graphene) with a thickness of 100 m. The sample surface was cleaved with an adhesive tape prior to measurements. SHG arises from a non-zero second-order susceptibility in materials without inversion symmetry [32, 33, 34]. In WTe2, the Td phase has no inversion symmetry, whereas the 1T′ phase does. Consequently, the Td phase exhibits SHG, while the 1T′ does not [21].
The optical setup for the TR-SHG measurements with a reflective geometry is shown in Fig. 1(a). For SHG measurements, an 800 nm (1.55 eV), 40-fs pulse at 100 kHz from the regenerative amplifier system (RegA9040) was used as input to the optical parametric amplifier (OPA9850) to generate a signal at a wavelength of 1230 nm for a probe to induce SHG. The residual 800 nm (70-fs) pulse was used as a pump that excites the sample. In the double-pulse TR-SHG experiment, the pump beam was split into two by a Michelson interferometer, and the time interval between the first and second pulses () was controlled by a linear stage with 10 m resolution [20]. The ratio of the fluences of the two pump pulses was set to approximately 1:1. In TR-SHG measurements, the pump and probe pulses were focused onto the sample normally and at a 45∘ incident angle, respectively, with a diameter of 60 m and 100 m. The pump fluence of a single pulse was varied from 1 to 6 mJ/cm2 to prevent sample damage, while the probe fluence was fixed at 3.2 mJ/cm2. The reflected probe (1230 nm) and SHG (615 nm) beams were directed to short-pass and band-pass filters, and only the SHG signal was delivered to fiber-coupling optics. The SHG signal was then delivered to a Si-photodiode through a multi-mode fiber (core diameter 400 m, NA=0.39). To obtain a time dependence of the SHG signal, the time delay () between the pump and probe pulses is modulated at 9.5 Hz by a shaker. All measurements were performed under ambient conditions at room temperature.
Before TR-SHG measurements, static SHG measurements were performed to clarify the SHG response from Td-WTe2. Figure 1(b) shows the spectrum observed for the incoming probe photon (1230 nm) and the resulting SHG signal (615 nm) emitted from the sample. Figure 1(c) shows the dependence of the polarization angle on the 1230 nm light pulses measured without pump light to investigate the response of SHG to crystal orientation. SHG intensity reaches its maximum for the -axis polarization and minimum for the -axis polarization. For TR-SHG measurements, the polarization was set to the -axis to maximize the SHG intensity. Here, the intense probe (3.2 ) was used to obtain a sufficient signal-to-noise ratio due to the weak second harmonic effect. It is noted that the intense probe does not influence PIPT because the fluence was selected within the region where the intensity of SHG was proportional to the square of the intensity of the incident probe , i.e., up to 4 mJ/cm2 (see details in Supplemental Material (SM))[35], as expected for a second-order nonlinear process. If the probe contributes to the PIPT, the SHG should decrease following the lattice symmetry changes [36, 22]. However, this is not the case here. Furthermore, according to the absorption coefficient obtained from our spectroscopic ellipsometry measurement, the effective photogenerated carrier density can be estimated to be 4.3 times larger for the pump than for the probe at the same fluence level (see SM) [35]. This additional measurement further supports the conclusion that the intense probe does not influence PIPT.
To explore the contribution of electronic excitation and shear phonon oscillation in structural change, we performed coherent control of the interlayer shear phonon. Figure 2 shows the time-domain signal of the SHG intensity change () obtained under double-pulse excitation together with the case of single-pulse excitation at 1.5 mJ/cm2 (see the inset). As seen in the inset, the SHG intensity decreases just after photoexcitation (2 ps) and then exhibits a long-lived coherent oscillation for tens of picoseconds. The frequency of the long-lived coherent oscillation was found to be 0.24 THz, matching the interlayer shear phonon assigned as the mode [21, 37, 29]. The time interval between the first and second pulses is 1.5 (= 6.25 ps) and 2 (= 8.33 ps) of the interlayer shear phonon at the identical total pump fluence of 4.6 mJ/cm2, where = 4.17 ps is the period of the interlayer shear phonon.
The decrease in SHG intensity indicates that the lattice symmetry changes from the Td toward 1T′ phases [21]. When the value of is an integral multiple of the phonon period, the oscillation amplitude of the coherent phonon generated by the first pulse is enhanced by the second pulse [11]. In contrast, when it is a half-integer multiple of the phonon period, it is canceled by the second pulse. As expected, the shear phonon was canceled when =1.5 and enhanced when =2. From Fig. 2, the only difference between the two cases is whether or not the oscillation of shear phonon, and no significant difference in the magnitude of was observed after 10 ps. Therefore, in the symmetry change from the Td toward 1T′ phases under the near-infrared (NIR) photon pump (800 nm), the effect of electronic excitation is a more dominant factor than the displacement of the shear phonon. In other words, the excitation of the coherent shear phonon can be independent of the formation of the metastable phase. Although decoupling of the shear phonon from the symmetry change from the Td toward 1T′ phases was suggested by suppressing the shear phonon with lattice defects [37], our results indicate more direct evidence by using coherent control of the phonon amplitude.
To gain more insight into the effects of coherent control of the phonon amplitude on the possible lattice symmetry change from the Td toward 1T′ phases, we present time-domain SHG signals for both constructive ( = 8.33 ps) and destructive ( = 6.25 ps) excitation at the total pump fluence from 1.6 (1.8) mJ/cm2 to 10.6 mJ/cm2 in Fig. 3. The amplitude of the shear phonon is enhanced by constructive excitation, as shown in Fig. 3(a), whereas it is canceled by destructive excitation as shown in Fig. 3(b). In both cases, the SHG intensity decreases as the total pump fluence increases.
To estimate the magnitude of for single- and double-pulse excitation, the time-domain data were fit separately using an exponentially decaying function , where is the Heaviside function convoluted with Gaussian to account for the finite time resolution, is the magnitude, and is the relaxation time constant. Note that the maximum normalized change of observed at the highest total fluence of 12 mJ/cm2 was 90%, which is comparable to that followed by the NIR pump pulse (2.1 m) with 10 MV/cm [21].
Figure 4 shows the pump fluence dependence of obtained from the fitting in Fig. 3. As seen in Fig. 4, higher pump fluence leads to a more significant decrease in the SHG intensity for both single- and double-pulse excitation. Moreover, even in the fluence dependence, the SHG intensity is significantly lower for double-pulse excitation than for single-pulse excitation. This suggests that saturable absorption was suppressed by multiple-pulse excitation [38, 25, 15] and/or that the structure is more likely to change if a second pump pulse is applied after the first has already modified the structure [20]. Although there has been a Landau theory to account for the temperature dependence of the SHG intensity [39], to our knowledge, an appropriate model for the fluence dependence will currently not be available under double-pulse excitation. Thus, the fluence dependence can be adapted to two linear functions with different slopes, implying a change of the potential energy surface [40, 1], e.g., a shift of the potential energy minimum toward the 1T′ phase, under PIPT. In this case, the critical fluence will be higher for double-pulse excitation (6 mJ/cm2) than for single-pulse excitation (4 mJ/cm2). More experimental and theoretical studies will be required to fully understand the fluence dependence found in Fig. 4.
To evaluate the impact of thermal effects during the phase transition from the Weyl semimetal Td to semimetal 1T′ phases, we have calculated the electron and lattice temperatures using the two-temperature model (TTM) [41] under double-pulse excitation as well as incorporating possible cumulative effects [42]. TTM, which applies to zero-gap metallic systems, can be used to examine the effect of elevated lattice temperatures. The TTM results for the two typical pump fluences in our experiment (7.0 mJ/cm2 and 9.2 mJ/cm2) are shown in SM[35]. Although the TTM analysis can only be used as a crude estimate, after the arrival of the second pulse, the lattice temperature () exceeds the transition temperature ( = 565 K[43] or 613 K [44]) at the time delay of 10 ps for 9.2 mJ/cm2. On the other hand, is below for 7.0 mJ/cm2. Interestingly, this critical fluence (7.0 mJ/cm2) is close to that observed for the depletion of the SHG intensity in Fig. 4, implying that the flatter part for 6 mJ/cm2 would need to be further addressed focusing on, in particular, early-time dynamics (first several picoseconds) where the shear phonon displacements facilitate the symmetry switch. Note that the actual temperature increase estimated may slightly differ from the TTM results because the TTM maintains homogeneous absorption, no multi-photon absorption, both of which could occur under high-density photoexcitation, this is the case here. Nevertheless, we can conclude that the lattice symmetry switching from the Weyl semimetallic Td to the semimetallic 1T′ phases is not promoted by the lattice temperature rise at the total pulse fluence of 7 mJ/cm2.
We argue that the lattice symmetry switch from the Td to the 1T′ phase is driven either by interband or intraband electronic excitation. The former promotes electrons from bonding to anti-bonding states, increasing the charge carrier density . In contrast, the latter intraband transition results in an increase in the electron temperature (or electron acceleration) without changing and can preferably be achieved by irradiating far-infrared or THz pulses [21]. Both excitation processes can promote carrier (hole) doping and induce shear sliding of atomic layers toward the 1T′ phase [9]. Under NIR (800 nm) conditions, the photoexcited carriers in WTe2 are generated by interband transitions along and directions in the electronic band structure [45]. In fact, we measured the imaginary part () of the complex refractive index of WTe2 by spectroscopic ellipsometry, and found that the 800 nm ( 1.55 eV) pump light used in the present study was resonant with 1.56 eV (see details in SM)[35], which can generate the photoexcited carrier via the interband transition. Thus, the increase in charge carrier density, e.g., 1.26 1021/cm3 (at 1.6 mJ/cm2) due to interband transitions plays a dominant role in the symmetry switch.
This study demonstrates that coherent control of the shear phonon amplitude can suppress its effect, isolating the impact of electronic excitation. We find that inversion-symmetry changes in the crystal structure mainly result from electronic excitation, i.e., incoherent PIPT [1]. Although previous attempts have used coherent shear phonons to induce structural transitions, we successfully decouple shear phonon motion from the Td-to-1T′ phase switch in WTe2 through destructive excitation. Since the TR-SHG technique is not a method that directly measures lattice displacement, in particular, atomic motions during this early time window (several picoseconds), it will be important to examine the present results using direct methods such as ultrafast electron diffraction [17, 21] and time-resolved x-ray diffraction [46, 47], which can measure the lattice displacement associated with the shear phonon. Even though, our findings pave the way for the investigation of phase transitions in other 2D materials, such as MoTe2, using constructive and destructive excitation of coherent shear phonons under electronic excitation.
In conclusion, we demonstrate coherent control of coherent shear phonons in the time-domain using pairs of femtosecond laser pulses to examine the role of interlayer shear phonons on the lattice symmetry switching in solids. Combining double-pulse excitation and time-resolved SHG techniques in WTe2, we reveal that the Td-to-1T′ symmetry switch is independent of the coherent shear phonon motion. Furthermore, on the basis of spectroscopic ellipsometry measurements, we establish that electronic excitation, particularly interband transitions, plays a central role in the phase transition, reinforcing its classification as incoherent PIPT. These results provide a framework for understanding the role of shear phonons in structural changes in 2D materials through coherent control of interlayer phonons.
Acknowledgement
This work was supported by JSPS KAKENHI (Grant Numbers. 25KJ0687, 25K17332, 23K22422 and 22KJ0352) and CREST, JST (Grant Number JPMJCR1875). T.F. acknowledges the support of the Sasakawa Scientific Research Grant from the Japan Science Society. Y.M. acknowledges the support from JST SPRING, Japan Grant Number JPMJSP2124. M.H. acknowledges the support from the DGIST R&D program (24-KUJoint-06). We thank Sathvik A. Iyengar and Gyan Prakash for their critical reading of the manuscript and insightful feedback, and Masaki Hada and the Organization for Open Facility Initiatives, University of Tsukuba, for the spectroscopic ellipsometry measurements.